# Optimal Computation in Anonymous Dynamic Networks

### v8

@@ 1.1 Background
Problem classes. Assume that each agent is assigned an input at the beginning of the computation, i.e., at round 00. The Input Frequency function is the function that returns the percentage of agents that have each input value. Being able to compute the Input Frequency function allows a system to compute a wide class of functions called frequency-based with no loss in performance [35]. Thus, we say that the Input Frequency function is complete for this class of functions (refer to Section 2.2). The most prominent representative of this class of functions is given by the aforementioned Average Consensus problem, since the mean of a multiset of numeric input values is a frequency-based function (the percentage of agents that have each input can be used as weight to compute the mean of all inputs).
@@ 1.3 Our Contributions
$\tau$ and $N\geq n$ known $\tau(n+N-2)+1$ $\tau(2n-4)$
$d$ known $\tau(n-1)+d+1\leq dn+1$
[…]
@@
Table 1: Summary of the results in this paper. The variable $n$ indicates the number of agents in the system, $\ell$ is the number of leaders, $N$ is an upper bound on $n$, $\tau$ is the dynamic disconnectivity of the network (or an upper bound thereof), and $d$ is its dynamic diameter (or an upper bound thereof). If $\ell=0$, only the frequency-based functions are computable; if $\ell\geq 1$, the multiset-based functions are also computable.
@@
We also provide lower bounds that asymptotically match our algorithms’ running times, assuming that the number of leaders $\ell$ is a constant (which is a realistic assumption in most applications). Table 1 summarizes our lower bounds, as well. We point out that our lower bounds of roughly $2\tau n$ rounds are the first non-trivial lower bounds for simple undirected anonymous dynamic networks (i.e., better than $n-1$).
@@ 2.1 Model of Computation
Input and internal states. Each agent $p_{i}$ starts with an input $\lambda(p_{i})$, which is assigned to it at round 00. It also has an internal state, which is initially determined by $\lambda(p_{i})$. At each round $t\geq 1$, every agent composes a message (depending on its internal state) and broadcasts it to its neighbors in $G_{t}$ through all its incident links. [^1] In order to model wireless radio communication, it is natural to assume that each agent in a dynamic network broadcasts its messages to all its neighbors (a message is received by anyone within communication range). The network’s anonymity prevents agents from specifying single destinations. By the end of round $t$, each agent reads all messages coming from its neighbors and updates its internal state according to a local algorithm $\mathcal{A}$. That is, $\mathcal{A}$ takes as input the current internal state of the agent and the multiset of incoming messages, and returns the new internal state of the agent. Note that $\mathcal{A}$ is deterministic and is the same for all agents.
@@ 2.3 Disconnected Networks
A limitation of Proposition 2.4 is that, unlike Proposition 2.3, it requires the presence of self-loops in the network. However, we can remove this requirement and prove a similar impossibility result for simple networks, at the cost of slightly narrowing the class of ”non-trivial functions” that cannot be computed.
@@
Sufficiency of $\tau=1$. The following result will be used repeatedly to reduce computations on $\tau$-union connected networks to computations on 1-union-connected networks. Essentially, it states that any algorithm for 1-union-connected networks can be extended to $\tau$-union connected networks (where $\tau$ is known to all agents) at the cost of a factor $\tau$ in the running time, which is optimal.
@@
We chose to predominantly use $\tau$, as opposed to $d$, to measure the running times of our algorithms for several reasons. Firstly, $\tau$ is well defined (i.e., finite) if and only if $d$ is; however, $\tau$ has a simpler definition, and is arguably easier to directly estimate or enforce in a real network. Secondly, Proposition 2.5, as well as all of our theorems, remain valid if we replace $\tau$ with $d$; nonetheless, stating the running times of our algorithms in terms of $\tau$ is preferable, because $\tau\leq d$.
@@ 3.1 Definition of History Tree
Indistinguishable agents. In an anonymous network, agents can only be distinguished by their inputs or by the multisets of messages they have received thus far. This leads to an inductive definition of indistinguishability: two agents are indistinguishable at the end of round 00 if and only if they have the same input. At the end of round $t\geq 1$, two agents $p$ and $q$ are indistinguishable if and only if they were indistinguishable at the end of round $t-1$ and, for every equivalence class $A$ of agents that were indistinguishable at the end of round $t-1$, both $p$ and $q$ receive an equal number of (identical) messages from agents in $A$ at round $t$.
@@
In addition, each node of $L_{0}$ has a label denoting the input of the agents it represents (by definition of indistinguishability at round 00, each node of $L_{0}$ represents all agents with a certain input). Nodes not in $L_{0}$ do not have any labels.
@@ 3.2 Construction of Vistas
Constructing the vista at round 00 is simple, because the agent just has to read its own input and create a root $r$ with a single child bearing a label equal to its input.
@@
If $t=0$, then $\mathcal{V}(p,t)$ consists of a single node $h(p,t)$ in $L_{0}$ attached to the root $r$. Since each node of $L_{0}$ is labeled as the input of the agents it represents, $\mathcal{F}_{\mathcal{A}}$ can extract the input of $p$ from $h(p,t)$ and use it to compute its internal state $\sigma(p,t)$, which at round 00 is just a function of the input.
@@
It is easy to infer $h(p,t)$ from $\mathcal{V}(p,t)$, because it is the unique node of maximum depth. Thus, the node $h(p,t-1)$ is the parent of $h(p,t)$, which is in $\mathcal{V}(p,t)$ because it is connected to $h(p,t)$ by a black edge. It follows that $\mathcal{F}_{\mathcal{A}}$ can extract from $\mathcal{V}(p,t)$ the vista of $p$ at round $t-1$ and, by the inductive hypothesis, it can compute $\sigma(p,t-1)$.
Also, for each agent $q$ that sends messages to $p$ at round $t$, the node $h(q,t-1)$ is in $\mathcal{V}(p,t)$, because it is connected to $h(p,t)$ by a red edge with a certain multiplicity $m\geq 1$. Hence, $\mathcal{F}_{\mathcal{A}}$ can extract $m$ from $\mathcal{V}(p,t)$, as well as the vista of $q$ at round $t-1$, and compute $\sigma(q,t-1)$ by the inductive hypothesis. From the internal state $\sigma(q,t-1)$, the message received by $p$ from $q$ at round $t$ can also be computed.
@@ 3.3 Related Concepts
Linear bounds. The $n^{2}$ upper bound was later improved by Norris, who showed that truncating views at depth $n-1$ is sufficient and may be necessary in some networks. That is, if two agents have isomorphic views up to depth $n-1$, they have isomorphic views at any depth [54].
@@
Broadcasting speed. As a consequence of the above, broadcasting a message, i.e., forwarding a message from a single agent to the whole dynamic network, takes at most $n-1$ rounds. Therefore, the dynamic diameter $d$, defined as the maximum number of rounds it takes to forward a message from an agent to any other agent at any point in time, satisfies $d\leq n-1$ in any 1-union-connected dynamic network.
@@
Let $P$ be a set of agents in a 1-union-connected dynamic network of size $n$, such that $11\leq|P|\leq n-$ 1, and let $t\geq 0$. Then, at every round $t^{\prime}\geq t+|P|$, in the history of every agent there is a node at level $L_{t}$ representing at least one agent not in $P$.
@@
Let $Q$ be the complement of $P$ (note that $Q$ is not empty), and let $Q_{t+i}$ be the set of agents represented by the nodes in $L_{t+i}$ whose vista contains a node in $L_{t}$ representing at least one agent in $Q$. We will prove by induction that $|Q_{t+i}|\geq|Q|+i$ for all $0\leq i\leq|P|$. The base case holds because $Q_{t}=Q$. The induction step is implied by $Q_{t+i}\subsetneq Q_{t+i+1}$, which holds for all $0\leq i<|P|$ as long as $|Q_{t+i}|<n$. Indeed, because $G_{t+i+1}$ is connected, it must contain a link between an agent $p\in Q_{t+i}$ and an agent $q\notin Q_{t+i}$. Thus, the history of $q$ at round $t+i+1$ contains the history of $p$ at round $t+i$, and so $Q_{t+i}\subsetneq Q_{t+i}\cup\{q\}\subseteq Q_{t+i+1}$.
Now, plugging $i:=|P|$, we get $|Q_{t+|P|}|=n$. In other words, the history of each node in $L_{t+|P|}$ (and hence in subsequent levels) contains a node at level $L_{t}$ representing an agent in $Q$. ∎
@@
Let $v\in L_{t}$, and let $P$ be the set of agents not represented by $v$. If $P$ is empty, then all nodes in $L_{t^{\prime}}$ are descendants of $v$, and have $v$ in their vista. Otherwise, $11\leq|P|\leq n-$ 1, and Lemma 3.3 implies that $v$ is in the vista of all nodes in $L_{t^{\prime}}$. ∎
@@ 4.1 Basic Technique
Our basic technique makes use of the procedure in Listing 1, which will be a subroutine of all algorithms presented in this section. The purpose of this procedure is to construct a homogeneous system of $k-1$ independent linear equations involving the anonymities of all the $k$ nodes in a level of an agent’s vista (recall that a linear system is homogeneous if all its constant terms are zero). We will first give some definitions.
@@
Main subroutine. Intuitively, the procedure in Listing 1 searches for a long-enough sequence of levels in the given vista $\mathcal{V}$, say from $L_{s}$ to $L_{t}$, where all nodes are non-branching. That is, the nodes in $L_{s}\cup L_{s+1}\cup\dots\cup L_{t}$ can be partitioned into $k=|L_{s}|=|L_{t}|$ strands. Then the procedure searches for pairs of exposed strands, each of which yields a linear equation involving the anonymities of some nodes of $L_{t}$, until it obtains $k-1$ linearly independent equations. The reason why we have to consider strands spanning several levels of the history tree (as opposed to looking at a single level) is that the dynamic disconnectivity $\tau$ is not known, and thus Proposition 2.5 cannot be applied directly. Note that the search may fail (in which case Listing 1 returns $t=-1$) or it may produce incorrect equations. The following lemma specifies sufficient conditions for Listing 1 to return a correct and non-trivial system of equations for some $t\geq 0$.
@@ 8 Else
9 Let $k=|L_{s}|=|L_{t}|$ and let $u_{i}$ be the $i$th node in $L_{t}$
10 Let $P_{i}$ be the strand starting in $L_{s}$ and ending in $u_{i}\in L_{t}$
11 Let $\mathcal{P}=\{P_{1},P_{2},\dots,P_{k}\}$
@@ 20 Return $(t,S)$
    $t^{\prime}\geq\tau(2n-c-1)+1$, where $c$ is the number of distinct inputs held by agents at round 00.
Then, $0\leq t\leq\tau(n-c)$, and $S$ is a homogeneous system of $k-1$ independent linear equations (with integer coefficients) in $k=|L_{t}|$ variables $x_{1}$, $x_{2}$, …, $x_{k}$. Moreover, $S$ is satisfied by assigning to $x_{i}$ the anonymity of the $i$th node of $L_{t}$, for all $1\leq i\leq k$.
@@
If $\tau=1$, it takes at most $n-1$ rounds for information to travel from an agent to any other agent, due to Corollary 3.4. In general, if $\tau\geq 1$, it takes at most $\tau(n-1)$ rounds, by Proposition 2.5. Therefore, since $\mathcal{V}$ is a vista taken at round $t^{\prime}$, all levels of $\mathcal{V}$ up to $L_{t^{\prime}-\tau(n-1)}$ are complete (recall that a level of a vista is complete if it has no missing nodes).
@@
Assume Condition 2 first. Since $t^{\prime}\geq\tau(2n-c-1)+1$, all levels of $\mathcal{V}$ up to $L_{\tau(n-c)+1}$ are complete. Thus, level $L_{0}$ is complete, and therefore it has exactly $c$ nodes, because two agents are distinguishable at round 00 if and only if they have distinct inputs. Since the sum of the anonymities of these $c$ nodes is $n$, there must be fewer than $n-c$ branching nodes in $\mathcal{V}$ (excluding the root). Hence, the levels up to $L_{\tau(n-c)}$ contain an interval of at least $\tau$ consecutive levels, say from $L_{r}$ to $L_{r+\tau-1}$, where all nodes are non-branching and can be partitioned into $|L_{r}|=|L_{r+\tau-1}|$ strands $P_{1}$, $P_{2}$, …, $P_{|L_{r}|}$.
@@
Note that a link between two agents at any round $r^{\prime}$ in the interval $[r+1,r+\tau]$ determines a pair of exposed nodes in $L_{r^{\prime}-1}$. Thus, by definition of $\tau$-union-connected network, the graph of exposed strands between $L_{r}$ and $L_{r+\tau-1}$ (constructed as $G$ in Line 12) is connected. It follows that the execution of Listing 1 terminates at Line 20 (as opposed to Line 6) whenever $t\geq r+\tau-1$. Thus, the procedure returns a pair $(t,S)$ with $0\leq t\leq r+\tau-1\leq\tau(n-c)$. In particular, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete.
@@
Now assume Condition 1. Since $t^{\prime}\geq t+\tau(n-1)+1$, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete in this case, as well. Since $t\geq 0$ by assumption, the execution of Listing 1 terminates at Line 20. The termination condition is met when long-enough strands are found; as proved above, this event must occur while $t\leq\tau(n-c)$.
@@
We have proved that, in both cases, the inequalities $0\leq t\leq\tau(n-c)$ hold, and all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete. Let us now examine the linear system $S$. Observe that $S$ is homogeneous because it consists of homogeneous linear equations (cf. Line 19). Also, since the spanning tree $G^{\prime}$ constructed at Line 14 has $k-1$ edges, $S$ contains $k-1$ equations. We will prove that they are linearly independent by induction on $k$. If $k=1$, there is nothing to prove. Otherwise, let $P_{i}$ be a leaf of $G^{\prime}$, and let $\{P_{i},P_{j}\}$ be its incident edge. Then, $S$ contains an equation $Q$ of the form $m_{1}x_{i}=m_{2}x_{j}$ with $m_{1}m_{2}\neq 0$. Let $S^{\prime}$ be the system obtained by removing $Q$ from $S$; equivalently, $S^{\prime}$ corresponds to the tree obtained by removing the leaf $P_{i}$ from $G^{\prime}$. By the inductive hypothesis, no linear combination of equations in $S^{\prime}$ yields $0=0$. On the other hand, if $Q$ is involved in a linear combination with a non-zero coefficient, then the variable $x_{i}$ cannot vanish, because it only appears in $Q$. Therefore, the equations in $S$ are independent.
@@
    $t^{\prime}\geq\tau(n-c)+d+1$, where $c$ is the number of distinct inputs held by agents at round 00,

### v9

@@ 1.1 Background
Problem classes. Assume that each agent is assigned an input at the beginning of the computation, i.e., at round 0. The Input Frequency function is the function that returns the percentage of agents that have each input value. Being able to compute the Input Frequency function allows a system to compute a wide class of functions called frequency-based with no loss in performance [35]. Thus, we say that the Input Frequency function is complete for this class of functions (refer to Section 2.2). The most prominent representative of this class of functions is given by the aforementioned Average Consensus problem, since the mean of a multiset of numeric input values is a frequency-based function (the percentage of agents that have each input can be used as weight to compute the mean of all inputs).
@@ 1.3 Our Contributions
$\tau$ and $N\geq n$ known $\tau(n+N-2)$ $\tau(2n-4)$
$d$ known $\tau(n-1)+d\leq\tau(2n-2)$ $\tau(2n-4)$
Multiset-based (e.g., Counting) $\ell\geq 1$ $\ell$ known $\tau(2n-2)$ $\tau(2n-\ell-5)$
@@
Table 1: Summary of the results in this paper. The variable $n$ indicates the number of agents in the system, $\ell$ is the number of leaders, NN is an upper bound on $n$, $\tau$ is the dynamic disconnectivity of the network (or an upper bound thereof), and $d$ is its dynamic diameter. If $\ell=0$, only the frequency-based functions are computable; if $\ell\geq 1$, the multiset-based functions are also computable. All the running times hold for multigraph networks, while all lower bounds hold even for simple networks.
@@
We also provide lower bounds that asymptotically match our algorithms’ running times, assuming that the number of leaders $\ell$ is a constant (which is a realistic assumption in most applications). Table 1 summarizes our lower bounds, as well. We point out that our lower bounds of roughly $2\tau n$ rounds are the first non-trivial lower bounds for simple undirected anonymous dynamic networks (i.e., better than n1n-1).
@@ 2.1 Model of Computation
Input and internal states. Each agent $p_{i}$ starts with an input $\lambda(p_{i})$, which is assigned to it at round 0. It also has an internal state, which is initially determined by $\lambda(p_{i})$. At each round $t\geq 1$, every agent composes a message (depending on its internal state) and broadcasts it to its neighbors in $G_{t}$ through all its incident links. [^1] In order to model wireless radio communication, it is natural to assume that each agent in a dynamic network broadcasts its messages to all its neighbors (a message is received by anyone within communication range). The network’s anonymity prevents agents from specifying single destinations. By the end of round $t$, each agent reads all messages coming from its neighbors and updates its internal state according to a local algorithm $\mathcal{A}$. That is, $\mathcal{A}$ takes as input the current internal state of the agent and the multiset of incoming messages, and returns the new internal state of the agent. Note that $\mathcal{A}$ is deterministic and is the same for all agents.
@@ 2.3 Disconnected Networks
A limitation of Proposition 2.4 is that, unlike Proposition 2.3, it requires the presence of self-loops in the network. However, we can remove this requirement and prove a similar impossibility result for simple networks, at the cost of slightly narrowing the class of “non-trivial functions” that cannot be computed.
@@
Sufficiency of $\tau=1$. The following result will be used repeatedly to reduce computations on $\tau$-union-connected networks to computations on 1-union-connected networks. Essentially, it states that any algorithm for 1-union-connected networks can be extended to $\tau$-union-connected networks (where $\tau$ is known to all agents) at the cost of a factor $\tau$ in the running time, which is optimal.
@@
We chose to predominantly use $\tau$, as opposed to $d$, to measure the running times of our algorithms for several reasons. Firstly, $\tau$ is well defined (i.e., finite) if and only if $d$ is; however, $\tau$ has a simpler definition, and is arguably easier to directly estimate or enforce in a real network. Secondly, Proposition 2.5, as well as all of our theorems in Sections 4 and 5, remain valid if we replace $\tau$ with $d$; nonetheless, stating the running times of our algorithms in terms of $\tau$ is preferable, because $\tau\leq d$.
@@ 3.1 Definition of History Tree
Indistinguishable agents. In an anonymous network, agents can only be distinguished by their inputs or by the multisets of messages they have received thus far. This leads to an inductive definition of indistinguishability: two agents are indistinguishable at the end of round 0 if and only if they have the same input. At the end of round $t\geq 1$, two agents $p$ and $q$ are indistinguishable if and only if they were indistinguishable at the end of round t1t-1 and, for every equivalence class $A$ of agents that were indistinguishable at the end of round t1t-1, both $p$ and $q$ receive an equal number of (identical) messages from agents in $A$ at round $t$.
@@
In addition, each node of $L_{0}$ has a label denoting the input of the agents it represents (by definition of indistinguishability at round 0, each node of $L_{0}$ represents all agents with a certain input). Nodes not in $L_{0}$ do not have any labels.
@@ 3.2 Construction of Vistas
Constructing the vista at round 0 is simple, because the agent just has to read its own input and create a root $r$ with a single child bearing a label equal to its input.
@@
If $t=0$, then $\mathcal{V}(p,t)$ consists of a single node $h(p,t)$ in $L_{0}$ attached to the root $r$. Since each node of $L_{0}$ is labeled as the input of the agents it represents, $\mathcal{F}_{\mathcal{A}}$ can extract the input of $p$ from $h(p,t)$ and use it to compute its internal state $\sigma(p,t)$, which at round 0 is just a function of the input.
@@
It is easy to infer $h(p,t)$ from $\mathcal{V}(p,t)$, because it is the unique node of maximum depth. Thus, the node h(p,t1)h(p,t-1) is the parent of $h(p,t)$, which is in $\mathcal{V}(p,t)$ because it is connected to $h(p,t)$ by a black edge. It follows that $\mathcal{F}_{\mathcal{A}}$ can extract from $\mathcal{V}(p,t)$ the vista of $p$ at round t1t-1 and, by the inductive hypothesis, it can compute $\sigma(p,t-1)$.
Also, for each agent $q$ that sends messages to $p$ at round $t$, the node h(q,t1)h(q,t-1) is in $\mathcal{V}(p,t)$, because it is connected to $h(p,t)$ by a red edge with a certain multiplicity $m\geq 1$. Hence, $\mathcal{F}_{\mathcal{A}}$ can extract $m$ from $\mathcal{V}(p,t)$, as well as the vista of $q$ at round t1t-1, and compute $\sigma(q,t-1)$ by the inductive hypothesis. From the internal state $\sigma(q,t-1)$, the message received by $p$ from $q$ at round $t$ can also be computed.
@@ 3.3 Related Concepts
Linear bounds. The $n^{2}$ upper bound was later improved by Norris, who showed that truncating views at depth n1n-1 is sufficient and may be necessary in some networks. That is, if two agents have isomorphic views up to depth n1n-1, they have isomorphic views at any depth [54].
@@
Broadcasting speed. As a consequence of the above, broadcasting a message, i.e., forwarding a message from a single agent to the whole dynamic network, takes at most n1n-1 rounds. Therefore, the dynamic diameter $d$, defined as the maximum number of rounds it takes to forward a message from an agent to any other agent at any point in time, satisfies $d\leq n-1$ in any 1-union-connected dynamic network.
@@
Let $P$ be a set of agents in a 1-union-connected dynamic network of size $n$, such that $1\leq|P|\leq n-1$, and let $t\geq 0$. Then, at every round $t^{\prime}\geq t+|P|$, in the vista of every agent there is a node at level $L_{t}$ representing at least one agent not in $P$.
@@
Let $Q$ be the complement of $P$ (note that $Q$ is not empty), and let $Q_{t+i}$ be the set of agents represented by the nodes in $L_{t+i}$ whose vista contains a node in $L_{t}$ representing at least one agent in $Q$. We will prove by induction that $|Q_{t+i}|\geq|Q|+i$ for all $0\leq i\leq|P|$. The base case holds because $Q_{t}=Q$. The induction step is implied by $Q_{t+i}\subsetneq Q_{t+i+1}$, which holds for all $0\leq i<|P|$ as long as $|Q_{t+i}|<n$. Indeed, because $G_{t+i+1}$ is connected, it must contain a link between an agent $p\in Q_{t+i}$ and an agent $q\notin Q_{t+i}$. Thus, the vista of $q$ at round $t+i+1$ contains the vista of $p$ at round $t+i$, and so $Q_{t+i}\subsetneq Q_{t+i}\cup\{q\}\subseteq Q_{t+i+1}$.
Now, plugging $i:=|P|$, we get $|Q_{t+|P|}|=n$. In other words, the vista of each node in $L_{t+|P|}$ (and hence in subsequent levels) contains a node at level $L_{t}$ representing an agent in $Q$. ∎
@@
Let $v\in L_{t}$, and let $P$ be the set of agents not represented by $v$. If $P$ is empty, then all nodes in $L_{t^{\prime}}$ are descendants of $v$, and have $v$ in their vista. Otherwise, $1\leq|P|\leq n-1$, and Lemma 3.3 implies that $v$ is in the vista of all nodes in $L_{t^{\prime}}$. ∎
@@ 4.1 Basic Technique
Our basic technique makes use of the procedure in Listing 1, which will be a subroutine of all algorithms presented in this section. The purpose of this procedure is to construct a homogeneous system of k1k-1 independent linear equations involving the anonymities of all the $k$ nodes in a level of an agent’s vista (recall that a linear system is homogeneous if all its constant terms are zero). We will first give some definitions.
@@
Main subroutine. Intuitively, the procedure in Listing 1 searches for a long-enough sequence of levels in the given vista $\mathcal{V}$, say from $L_{s}$ to $L_{t}$, where all nodes are non-branching. That is, the nodes in $L_{s}\cup L_{s+1}\cup\dots\cup L_{t}$ can be partitioned into $k=|L_{s}|=|L_{t}|$ strands. Then the procedure searches for pairs of exposed strands, each of which yields a linear equation involving the anonymities of some nodes of $L_{t}$, until it obtains k1k-1 linearly independent equations. The reason why we have to consider strands spanning several levels of the history tree (as opposed to looking at a single level) is that the dynamic disconnectivity $\tau$ is not known, and thus Proposition 2.5 cannot be applied directly. Note that the search may fail (in which case Listing 1 returns t=1t=-1) or it may produce incorrect equations. The following lemma specifies sufficient conditions for Listing 1 to return a correct system of non-trivial equations for some $t\geq 0$.
@@ 8 Else
9 Let $k=|L_{s}|=|L_{t}|$ and let $u_{1}$, $u_{2}$, $\dots$, $u_{k}$ be the nodes in $L_{t}$
10 Let $\mathcal{P}=\{P_{1},P_{2},\dots,P_{k}\}$, where $P_{i}$ is the strand starting in $L_{s}$ and ending at $u_{i}\in L_{t}$
@@ 19 Return $(t,S)$
    $t^{\prime}\geq\tau(2n-c)$, where $c$ is the number of distinct inputs held by agents at round 0.
Then, $0\leq t\leq\tau(n-c+1)-1$, and $S$ is a homogeneous system of k1k-1 independent linear equations (with integer coefficients) in $k=|L_{t}|$ variables $x_{1}$, $x_{2}$, …, $x_{k}$. Moreover, $S$ is satisfied by assigning to $x_{i}$ the anonymity of the $i$th node of $L_{t}$, for all $1\leq i\leq k$.
@@
If $\tau=1$, it takes at most n1n-1 rounds for information to travel from an agent to any other agent, due to Corollary 3.4. In general, if $\tau\geq 1$, it takes at most $\tau(n-1)$ rounds, by Proposition 2.5. Therefore, since $\mathcal{V}$ is a vista taken at round $t^{\prime}$, all levels of $\mathcal{V}$ up to $L_{t^{\prime}-\tau(n-1)}$ are complete (recall that a level of a vista is complete if it has no missing nodes).
@@
Assume Condition 2 first. Since $t^{\prime}\geq\tau(2n-c)$, all levels of $\mathcal{V}$ up to $L_{\tau(n-c+1)}$ are complete. Thus, level $L_{0}$ is complete, and therefore it has exactly $c$ nodes, because two agents are distinguishable at round 0 if and only if they have distinct inputs. Since the sum of the anonymities of these $c$ nodes is $n$, there may be at most ncn-c branching nodes in $\mathcal{V}$ (excluding the root). Hence, the levels from $L_{0}$ up to $L_{\tau(n-c+1)-1}$ contain an interval of at least $\tau$ consecutive levels, say from $L_{r}$ to $L_{r+\tau-1}$, where all nodes are non-branching and can be partitioned into $k=|L_{r}|=|L_{r+\tau-1}|$ strands $P_{1}$, $P_{2}$, …, $P_{k}$ (Lines 9–10).
@@
Note that a link between two agents at any round $r^{\prime}$ in the interval $r+1,r+\tau]$ determines a pair of exposed nodes in $L_{r^{\prime}-1}$. Thus, by definition of $\tau$-union-connected network, the graph of exposed strands between $L_{r}$ and $L_{r+\tau-1}$ (constructed as $G$ in Line 11) is connected. It follows that the execution of Listing 1 terminates at Line 19 (as opposed to Line 6) whenever $t\geq r+\tau-1$. Thus, the procedure returns a pair $(t,S)$ with $0\leq t\leq r+\tau-1\leq\tau(n-c+1)-1$. In particular, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete, since $t+1\leq\tau(n-c+1)$.
@@
Now assume Condition 1. Since $t^{\prime}\geq(t+1)+\tau(n-1)$, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete in this case, as well. Since $t\geq 0$ by assumption, the execution of Listing 1 terminates at Line 19. The termination condition is met when long-enough strands are found; as proved above, this event must occur when $t\leq\tau(n-c+1)-1$.
@@
We have proved that, in both cases, the inequalities $0\leq t\leq\tau(n-c+1)-1$ hold, and all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete. Let us now examine the linear system $S$. Observe that $S$ is homogeneous because it consists of homogeneous linear equations (cf. Line 18). Also, since the spanning tree $G^{\prime}$ constructed at Line 13 has k1k-1 edges, $S$ contains k1k-1 equations. We will prove that they are linearly independent by induction on $k$. If $k=1$, there is nothing to prove. Otherwise, let $P_{i}$ be a leaf of $G^{\prime}$, and let $\{P_{i},P_{j}\}$ be its incident edge. Then, $S$ contains an equation $Q$ of the form $m_{1}x_{i}=m_{2}x_{j}$ with $m_{1}m_{2}\neq 0$. Let $S^{\prime}$ be the system obtained by removing $Q$ from $S$; equivalently, $S^{\prime}$ corresponds to the tree obtained by removing the leaf $P_{i}$ from $G^{\prime}$. By the inductive hypothesis, no linear combination of equations in $S^{\prime}$ yields $0=0$. On the other hand, if $Q$ is involved in a linear combination with a non-zero coefficient, then the variable $x_{i}$ cannot vanish, because it only appears in $Q$. Therefore, the equations in $S$ are independent.
@@
    $t^{\prime}\geq\tau(n-c+1)+d$, where $c$ is the number of distinct inputs held by agents at round 0,